OpenStudy (anonymous):

What polynomial has roots of -4, 2, and 5 ?

OpenStudy (anonymous):

(x+4) (x-2) (x-5)

OpenStudy (anonymous):

can you explain how you found it

OpenStudy (anonymous):

Also, I think my teacher wants it in an equation

OpenStudy (anonymous):

if x = -4 then x + 4 = 0 x = 2, then x - 2 = 0 x = 5, then x - 5 = 0 (x+4) (x-2) (x-5) you have foil them to get the polynomial

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I got x^3 - 3x^2 -18x + 40

OpenStudy (mathmale):

WM: sourwing is absolutely right. You'll need to multiply out (x+4)(x-2)(x-5) to obtain the polynomial. Write it as f(x) = x^3 + ? + ? + ? (four terms).

OpenStudy (mathmale):

Now you could check whether that is correct using synthetic division, if you know synth. div.: Start with x =2: 2 | 1 -3 -18 +40 2 ------------------------ 1

OpenStudy (mathmale):

Or you could do your multiplication in a different order and see whether you obtain the same poloynomial.

OpenStudy (anonymous):

ok thanks for your help

OpenStudy (mathmale):

You're welcome! But are you going to check your answer?